OpenStudy (anonymous):

The slope of a line is -1/5, and the y-intercept is 5. What is the equation of the line written in general form? A. x - 5y - 25 = 0 B. x - 5y - 5 = 0 C. x + 5y - 25 = 0

OpenStudy (anonymous):

i recommend first putting your line into slope-intercept form, so. y=-1/5x+5

OpenStudy (anonymous):

your first step is to multiply both sides of your equation by the LCD of your fractions. you only have one fraction in your problem so multiply both sides by 5

OpenStudy (anonymous):

so now you have -5y=x-25

OpenStudy (anonymous):

and now solve for you b or your y-intercept which in your case is -25.

OpenStudy (anonymous):

its depends on what you prefer but if i was you i would add the 25 and the 5y to either side that way all of your values are positive so 25 = x+5y

OpenStudy (anonymous):

Thank you, I got the answer right! (: @TechnoSoul